OpenStudy (anonymous):

This One Is v.v Ugh . 6x+2y=24 ; Solve For X .

OpenStudy (ghazi):

6x+2y=24 subtract 2y from both the sides 6x= 24-2y now divide by 6 on both the sides x=(24-2y)/6=2(12-y)/6=(12-y)/3

OpenStudy (anonymous):

All I Got Is .. 6x=24-2y

OpenStudy (anonymous):

& Now I Got 6x=12-y

OpenStudy (anonymous):

x=-2y ?

OpenStudy (ghazi):

after that divide by 6 on both the sides\[\frac{ 6x }{ 6 }=\frac{ 24-2y }{ 6 }=\frac{ 12-2y }{ 1 }=x\]

OpenStudy (anonymous):

x=2 ?

OpenStudy (ghazi):

x=12-2y/3

OpenStudy (ghazi):

\[x=\frac{ 12-2y }{ 3 }\]

OpenStudy (ghazi):

\[x=\frac{ 2(12-2y) }{ 6 }=\frac{ 12-2y }{ 3 }\]

OpenStudy (anonymous):

Okay . Thank You Soo Much By The Way .

OpenStudy (ghazi):

sorry it's x= 12-y/3 :) YW

OpenStudy (anonymous):

Ohh Okay (: